# SciFIBench: Benchmarking Large Multimodal Models for Scientific Figure Interpretation

### v1

@@ 4.2 Experimental Settings
Inference. For the closed-source models, inference was carried out via the OpenAI API666https://platform.openai.com/docs/api-reference/ or Vertex AI API777https://cloud.google.com/vertex-ai/. We use the HuggingFace Transformers library [53] and OpenCompass toolkit [10] to access the open-source models and conduct inference using NVIDIA A100 GPUs. With current pricing, evaluating GPT-4V on the SciFIBench gold set costs $15. For the open-source models, the typical gold set inference runtime using an A100 is 45 minutes (e.g., using Qwen-VL).
Hyperparameters. We select model hyperparameter settings that produce deterministic output to encourage reproducibility. For the open-source models, we utilise the greedy search decoding strategy, in which the most probable token is selected from the model vocabulary at each step, conditional on the preceding tokens i.e., . For the Gemini and Claude 3 models, we set the to 0 and to 1; for the GPT-4 models, we also set the to 0 and specify a random seed.
@@
Evaluation Metrics.
We utilise a single evaluation metric when determining performance: the proportion of questions answered correctly. The same metric is used across both tasks.
@@
Despite instruction in the prompt to constrain the format of the model answers to each question to just the target choice letter, e.g., ‘A’, most of the evaluated models did not consistently follow this, posing a challenge to automatic evaluation (string comparison). To overcome this, we used Gemini-Pro to initially parse the output and extract the answer letter or flag if no single answer was given.
@@ 4.3 Main Results
To gauge the abilities of frontier LMMs to interpret scientific figures, we evaluate a diverse set of LMMs and other baselines on our SciFIBench gold dataset, the results for which are displayed in Tab. 2. Note, our core analysis is in reference to results obtained on the adversarially generated question negatives (columns 2 and 4). We present our key findings as follows:
| Accuracy | ||||||||||||||
| Model | FigureCaption | CaptionFigure | ||||||||||||
| Closed-source LMMs | ||||||||||||||
| GPT-4V [38] | 69.4 | +29.8 | 99.2 | 58.4 | +38.0 | 96.4 | ||||||||
| GPT-4o [39] | 75.4 | +24.2 | 99.6 | 72.2 | +26.8 | 99.0 | ||||||||
| Gemini-Pro Vision [49] | 56.0 | +41.2 | 97.2 | 52.4 | +46.0 | 98.4 | ||||||||
| Gemini-Pro 1.5 [43] | 74.0 | +25.0 | 99.0 | 76.0 | +22.4 | 98.4 | ||||||||
| Claude 3 Haiku [5] | 52.6 | +36.4 | 89.0 | 43.8 | +34.6 | 78.4 | ||||||||
| Claude 3 Sonnet [5] | 53.4 | +33.0 | 86.4 | 58.4 | +31.6 | 90.0 | ||||||||
| Claude 3 Opus [5] | 59.8 | +27.0 | 88.2 | 49.2 | +32.0 | 81.2 | ||||||||
| Open-source LMMs | ||||||||||||||
| IDEFICS-9b-Instruct [26] | 20.6 | +4.4 | 25.0 | 20.2 | -3.0 | 17.2 | ||||||||
| IDEFICS-80b-Instruct [26] | 20.6 | +17.6 | 38.2 | 24.2 | +0.4 | 24.6 | ||||||||
| Qwen-VL-Chat [6] | 28.0 | +30.0 | 58.0 | 16.0 | +1.0 | 17.0 | ||||||||
| Emu2 [47] | 20.8 | +28.4 | 49.2 | - | - | - | ||||||||
| TransCore-M [41] | 51.0 | +28.2 | 79.2 | - | - | - | ||||||||
| InternLM-XComposer-7b [50] | 34.0 | +21.6 | 55.6 | - | - | - | ||||||||
| InternLM-XComposer2-7b [50] | 28.0 | +46.0 | 74.0 | - | - | - | ||||||||
| CogVLM-Chat [52] | 40.8 | +17.0 | 57.8 | - | - | - | ||||||||
| OmniLMM-3b [40] | 35.8 | +29.0 | 64.8 | - | - | - | ||||||||
| OmniLMM-12b [40] | 34.2 | +34.0 | 68.2 | - | - | - | ||||||||
| Yi-VL-6b [1] | 41.4 | +30.4 | 71.8 | - | - | - | ||||||||
| Yi-VL-34b [1] | 32.6 | +29.4 | 62.0 | - | - | - | ||||||||
| InstructBLIP-FlanT5-xl [11] | 35.8 | +22.2 | 58.0 | - | - | - | ||||||||
| InstructBLIP-FlanT5-xxl [11] | 36.2 | +20.4 | 56.6 | - | - | - | ||||||||
| InstructBLIP-Vicuna-7b [11] | 21.0 | -3.4 | 17.6 | - | - | - | ||||||||
| InstructBLIP-Vicuna-13b [11] | 22.2 | +5.2 | 27.4 | - | - | - | ||||||||
| Monkey-Chat [30] | 27.2 | +22.8 | 50.0 | - | - | - | ||||||||
| LLaVA-1.5-7b [31] | 32.8 | +27.8 | 60.6 | - | - | - | ||||||||
| LLaVA-1.5-13b [31] | 25.0 | +41.2 | 66.2 | - | - | - | ||||||||
| VLMs | ||||||||||||||
| CLIP ViT-H-14-378-quickgelu [20] | 41.8 | +50.6 | 92.4 | 42.6 | +53.4 | 96.0 | ||||||||
| MetaCLIP ViT-H-14-quickgelu [57] | 36.6 | +53.2 | 89.8 | 35.4 | +54.8 | 90.2 | ||||||||
| Google Multimodal Embedding [2] | 47.6 | +46.2 | 93.8 | 54.4 | +44.0 | 98.4 | ||||||||
| Human (25 questions per task*) | ||||||||||||||
| Human () | 86.48.24 | +10.7 | 100.0 | 78.48.24 | +22.7 | 100.0 | ||||||||
| GPT-4o | 72.0 | +28.0 | 100.0 | 76.0 | +24.0 | 100.0 | ||||||||
| Gemini-Pro 1.5 | 84.0 | +16.0 | 100.0 | 72.0 | +28.0 | 100.0 | ||||||||
| Claude 3 Opus | 72.0 | +8.0 | 80.0 | 56.0 | +32.0 | 88.0 | ||||||||
| GPT-4V | 68.0 | +28.0 | 96.0 | 56.0 | +40.0 | 96.0 | ||||||||
| CLIP ViT-H-14-378-quickgelu | 48.0 | +44.0 | 92.0 | 56.0 | +44.0 | 100.0 | ||||||||
| TransCore-M | 36.0 | +48.0 | 84.0 | - | - | - | ||||||||
@@
SciFIBench represents a difficult benchmark. The best-performing models, GPT-4o and Gemini-Pro 1.5, attain scores of 75.4% and 74.0% for the Figure Caption task, respectively, and 72.2% and 76.0% for the Caption Figure tasks, respectively. This shows that even at the frontier there is sufficient headroom for improvement.
@@
Among the weaker models, there is much more headroom, with the weakest models only just equalling or barely surpassing the chance score. Overall, there is a large spread of performance scores across the models, suggesting the benchmark has a good range of question difficulties.
@@
Closed-source models are noticeably better than open-source models. We observe a large performance gap between the closed-source and open-source models. Considering the Figure Caption task, there is a difference of 24.4% between the scores of the best closed and open-sourced models. Moreover, the best-performing open-source model, TransCore-M underperforms the worst-performing closed-source model, Claude 3 Haiku. This difference is more pronounced for the Caption Figure task.
@@
Caption Figure is more difficult than Figure Caption. Among the closed-source models, slightly higher overall scores are attained on the Figure Caption task. This observation holds for the open-source models, especially in the random negatives setting. Considering the human baseline, a noticeably lower score is attained on the Caption Figure task, suggesting it is easier for humans to distinguish fine-grained details in the text domain. The VLM baselines show no discernible difference in performance across the tasks, a possible reflection of their pretraining strategy of jointly aligning language and vision.
@@
Considering the models that we evaluate different checkpoint sizes (i.e., IDEFICS, OmniLMM, Yi, InstructBLIP-FlanT5, InstructBLIP-Vicuna, LLaVA-1.5, InternLM), we find that more often than not, the smaller model outperforms the larger checkpoint on the adversarially selected negatives, however, the opposite is true for the randomly selected negatives. Additionally, the difference in performance is more pronounced on the randomly selected negatives.
@@
Humans are a stronger baseline. The mean human baseline outperforms all the models, though does not achieve a perfect score, reflecting the challenging nature of this benchmark and the fact that the participants were not necessarily domain experts. As indicated by the standard deviation, a range of accuracy scores were recorded for each task, with some participants scoring equal or lower than the best LMMs. It is worth noting a slight caveat to the human performance is that the human verification part of the dataset curation process could have introduced bias toward questions that are ‘easier’ for humans to answer.
@@
In the Appendix, we include qualitative results for each task and examples of model output before automatic evaluation.
4.4 Performance Across Categories
We analyse the performance of a representative sample of models on the Figure Caption task across the 12 categories in the SciFIBench gold dataset. The results are displayed in Fig. 5 and comprehensive results for all models can be found in the Appendix. Although the different categories in our benchmark are all (at least partly) from the arXiv CS category, there is considerable variation in the style and type of figures within each category. This disparity is reflected in the difference in performance across categories: the average accuracy for the selected models differs by 22.5% between the best category (LG) and the worse category (AI). There is slight variation in ranking of the models across categories, but it remains fairly consistent.
4.5 Gold vs. Silver Datasets
@@ 4.4 Performance Across Categories
Due to its increased quality and feasibility of evaluation, our analysis is focused on the gold set, with our main intentions for the silver and bronze sets being downstream fine-tuning and few-shot evaluation. Here, we justify this decision and provide evidence that although our gold set is relatively small, it is sufficiently robust. We evaluate a subset of our models on both the gold and silver sets and plot their relative rankings in Fig. 6. For almost all models, the ranking is preserved across the datasets, and in the case where the rankings switch, the performance differential between the two models is small. A key purpose of a benchmark is to provide the rankings of models. This analysis shows that this indicator is largely sustained, suggesting there is little information to be gained by evaluating on an arbitrarily larger dataset. Moreover, we conduct bootstrapping to estimate the variance of model performance on the gold dataset (500 questions for each task). Concretely, for each task, we sample with replacement 500 times from the relevant question set and evaluate the performance of Gemini Pro Vision (middle-performing model capable of both tasks) on the sample. Repeating this process 100k times yields a mean accuracy and variance of (56.00, 0.05) and (52.40, 0.05) for the Figure Caption and Caption Figure tasks, respectively. This low variance provides further evidence that our gold dataset is sufficiently representative.

### v2

@@ 4.2 Experimental Settings
Inference. For the closed-source models, inference was carried out via the OpenAI API [67] or Vertex AI API [68]. We use Transformers [69] and OpenCompass toolkit [70] to access the open-source models and conduct inference using NVIDIA A100 GPUs. With current pricing, evaluating GPT-4V on SciFIBench costs $30. For the open-source models, the typical inference runtime using an A100 is 1 hour (e.g., using Qwen-VL).
Hyperparameters. We select model hyperparameters that produce deterministic output. For the open-source models, we utilise the greedy search decoding strategy, in which the most probable token is selected from the model vocabulary at each step, conditional on the preceding tokens i.e., . For the Gemini and Claude models, we set the to 0 and to 1; for the GPT-4 models, we also set the to 0.
@@
Despite instruction to constrain the format of the model answers to each question to just the target choice letter, e.g., ‘A’, most of the evaluated models did not consistently follow this, posing a challenge to automatic evaluation. To overcome this, we used Gemini-Pro to initially parse the output and extract the answer letter or flag if no single answer was given.
@@ 4.3 Main Results
To gauge the abilities of frontier LMMs to interpret scientific figures, we evaluate a diverse set of LMMs and other baselines on SciFIBench, the results for which are displayed in Tab. 2. Note, our core analysis is in reference to results obtained on the adversarially generated question negatives. We present our key findings as follows:
| CS | General | Overall | ||||||||||||||||||||||||
| Model | Fig.Cap. | Cap.Fig. | Fig.Cap. | Cap.Fig. | Fig.Cap. | Cap.Fig. | ||||||||||||||||||||
| Closed-source LMMs | ||||||||||||||||||||||||||
| GPT-4V [5] | 69.4 | +29.8 | 99.2 | 58.4 | +38.0 | 96.4 | - | - | - | - | ||||||||||||||||
| GPT-4 Turbo [52] | 68.0 | +30.6 | 98.6 | 60.6 | +36.8 | 97.4 | 62.8 | 55.2 | 65.4 | 57.9 | ||||||||||||||||
| GPT-4o [31] | 75.4 | +24.2 | 99.6 | 72.2 | +26.8 | 99.0 | 72.2 | 58.6 | 73.8 | 65.4 | ||||||||||||||||
| Gemini Pro Vision [7] | 56.0 | +41.2 | 97.2 | 52.4 | +46.0 | 98.4 | 50.6 | 39.6 | 53.3 | 46.0 | ||||||||||||||||
| Gemini 1.5 Pro [32] | 74.0 | +25.0 | 99.0 | 76.0 | +22.4 | 98.4 | 65.2 | 56.2 | 69.6 | 66.1 | ||||||||||||||||
| Gemini 1.5 Flash [32] | 74.4 | +24.6 | 99.0 | 69.6 | +29.4 | 99.0 | 65.8 | 62.4 | 70.1 | 66.1 | ||||||||||||||||
| Claude 3 Haiku [53] | 52.6 | +36.4 | 89.0 | 43.8 | +34.6 | 78.4 | 52.6 | 33.0 | 52.6 | 38.4 | ||||||||||||||||
| Claude 3 Sonnet [53] | 53.4 | +33.0 | 86.4 | 58.4 | +31.6 | 90.0 | 53.6 | 55.0 | 53.5 | 56.7 | ||||||||||||||||
| Claude 3 Opus [53] | 59.8 | +27.0 | 88.2 | 49.2 | +32.0 | 81.2 | 50.8 | 47.4 | 55.3 | 48.3 | ||||||||||||||||
| Open-source LMMs | ||||||||||||||||||||||||||
| IDEFICS-9b-Instruct [54] | 20.6 | +4.4 | 25.0 | 20.2 | -3.0 | 17.2 | 17.6 | 12.6 | 19.1 | 16.4 | ||||||||||||||||
| IDEFICS-80b-Instruct [54] | 20.6 | +17.6 | 38.2 | 24.2 | +0.4 | 24.6 | 18.4 | 20.6 | 19.5 | 22.4 | ||||||||||||||||
| Qwen-VL-Chat [6] | 28.0 | +30.0 | 58.0 | 16.0 | +1.0 | 17.0 | 17.0 | 19.2 | 22.5 | 17.6 | ||||||||||||||||
| Emu2 [55] | 20.8 | +28.4 | 49.2 | - | - | - | 19.6 | - | 20.2 | - | ||||||||||||||||
| TransCore-M [56] | 51.0 | +28.2 | 79.2 | - | - | - | 27.4 | - | 39.2 | - | ||||||||||||||||
| InternLM-XComposer-7b [57] | 34.0 | +21.6 | 55.6 | - | - | - | 21.6 | - | 27.8 | - | ||||||||||||||||
| InternLM-XComposer2-7b [57] | 28.0 | +46.0 | 74.0 | - | - | - | 23.8 | - | 25.9 | - | ||||||||||||||||
| CogVLM-Chat [58] | 40.8 | +17.0 | 57.8 | - | - | - | 24.0 | - | 32.4 | - | ||||||||||||||||
| OmniLMM-3b [59] | 35.8 | +29.0 | 64.8 | - | - | - | 24.8 | - | 30.3 | - | ||||||||||||||||
| OmniLMM-12b [59] | 34.2 | +34.0 | 68.2 | - | - | - | 27.2 | - | 30.7 | - | ||||||||||||||||
| Yi-VL-6b [60] | 41.4 | +30.4 | 71.8 | - | - | - | 27.0 | - | 34.2 | - | ||||||||||||||||
| Yi-VL-34b [60] | 32.6 | +29.4 | 62.0 | - | - | - | 21.4 | - | 27.0 | - | ||||||||||||||||
| InstructBLIP-FlanT5-xl [61] | 35.8 | +22.2 | 58.0 | - | - | - | 19.0 | - | 27.4 | - | ||||||||||||||||
| InstructBLIP-FlanT5-xxl [61] | 36.2 | +20.4 | 56.6 | - | - | - | 26.8 | - | 31.5 | - | ||||||||||||||||
| InstructBLIP-Vicuna-7b [61] | 21.0 | -3.4 | 17.6 | - | - | - | 12.8 | - | 16.9 | - | ||||||||||||||||
| InstructBLIP-Vicuna-13b [61] | 22.2 | +5.2 | 27.4 | - | - | - | 15.6 | - | 18.9 | - | ||||||||||||||||
| Monkey-Chat [62] | 27.2 | +22.8 | 50.0 | - | - | - | 18.2 | - | 22.7 | - | ||||||||||||||||
| LLaVA-1.5-7b [63] | 32.8 | +27.8 | 60.6 | - | - | - | 22.8 | - | 27.8 | - | ||||||||||||||||
| LLaVA-1.5-13b [63] | 25.0 | +41.2 | 66.2 | - | - | - | 20.2 | - | 22.6 | - | ||||||||||||||||
| Text-only input | ||||||||||||||||||||||||||
| Gemini 1.5 Flash [32] | 48.0 | +33.4 | 81.4 | 39.2 | +34.4 | 73.6 | 51.0 | 35.8 | 49.5 | 37.5 | ||||||||||||||||
| VLMs | ||||||||||||||||||||||||||
| CLIP ViT-H-14-378-quickgelu [50] | 41.8 | +50.6 | 92.4 | 42.6 | +53.4 | 96.0 | 30.6 | 30.0 | 36.2 | 36.3 | ||||||||||||||||
| MetaCLIP ViT-H-14-quickgelu [65] | 36.6 | +53.2 | 89.8 | 35.4 | +54.8 | 90.2 | 24.2 | 25.2 | 30.4 | 30.3 | ||||||||||||||||
| Google Multimodal Embedding [66] | 47.6 | +46.2 | 93.8 | 54.4 | +44.0 | 98.4 | 28.2 | 28.4 | 37.9 | 41.4 | ||||||||||||||||
| Human (25 questions per task*) | ||||||||||||||||||||||||||
| Human () | 86.48.24 | +10.7 | 100.0 | 78.48.24 | +22.7 | 100.0 | - | - | - | - | ||||||||||||||||
| GPT-4o | 72.0 | +28.0 | 100.0 | 76.0 | +24.0 | 100.0 | - | - | - | - | ||||||||||||||||
| Gemini-Pro 1.5 | 84.0 | +16.0 | 100.0 | 72.0 | +28.0 | 100.0 | - | - | - | - | ||||||||||||||||
| CLIP ViT-H-14-378-quickgelu | 48.0 | +44.0 | 92.0 | 56.0 | +44.0 | 100.0 | - | - | - | - | ||||||||||||||||
| TransCore-M | 36.0 | +48.0 | 84.0 | - | - | - | - | - | - | - | ||||||||||||||||
@@
SciFIBench represents a difficult benchmark. The best-performing models, GPT-4o and Gemini 1.5 Flash, attain scores of 73.8% and 70.1% for the FigureCaption task, and 65.4% and 66.1% for the CaptionFigure tasks, respectively. This shows that even at the frontier there is room for improvement.
@@
Among the weaker models, there is much more headroom, with the weakest models only just equalling or surpassing the chance score. Overall, there is a large spread of performance scores across the models, suggesting the benchmark has a good range of question difficulties.
@@
Closed-source models are noticeably better than open-source models. Considering the Figure Caption task, there is a difference of 34.6% between the scores of the best closed and open-sourced models. Moreover, the best-performing open-source model, TransCore-M underperforms the worst closed-source model. This difference is more pronounced for the CaptionFigure task.
@@
CaptionFigure is more difficult than FigureCaption. Multi-image tasks are known to be challenging to LMMs [72], and slightly higher overall scores are attained on the FigureCaption task, especially in the random negatives setting. Considering the human baseline, a noticeably lower score is attained on the CaptionFigure task, suggesting it is easier for humans to distinguish fine-grained details in the text domain. The VLM baselines show no discernible difference in performance across the tasks, a possible reflection of their pretraining strategy of jointly aligning language and vision.
@@
Considering the models that we evaluate multiple checkpoint sizes (e.g., IDEFICS, OmniLMM, Yi, etc.), we find that more often than not, the smaller model outperforms the larger checkpoint on the adversarially selected negatives, however, the opposite is true for the randomly selected negatives. Additionally, the difference in performance is more pronounced on the randomly selected negatives.
@@
Humans are a stronger baseline. The mean human baseline outperforms all the models, though does not achieve a perfect score, reflecting the challenging nature of SciFIBench and the fact that the participants were not necessarily domain experts. As indicated by the standard deviation, a range of accuracy scores were recorded for each task, with some participants scoring equal or lower than the best LMMs. It is worth noting a caveat to the human performance is that the human verification part of the curation process could have introduced bias toward questions that are ‘easier’ for humans to answer.
OCR. As detection of fine-grained textual detail is a key component of scientific figure interpretation, it is not unexpected that an above-chance score can be attained in a text-only setting. However, given the significant difference in performance between the text-only and image settings, it is clear that interpretation of visual details is required to answer the majority of questions in our benchmark.
@@
In the Appendix, we include qualitative results for each task and examples of model output.
| k-shot | Similar | Random |
| examples | examples | |
| (µ±SD) | ||
| 0 | 56.8 | 56.8 |
| 1 | 57.6 | 55.8±1.2 |
| 2 | 61.2 | 59.4±0.6 |
| 3 | 59.2 | 58.4±1.0 |
| 4 | 58.0 | 58.3±1.0 |
| 5 | 58.0 | 58.0±1.4 |
4.4 Curated vs. noisy data
@@
Here, we provide evidence that although our benchmark is relatively small, it is sufficiently robust. We evaluate a subset of our models on both the curated and noisy CS question sets and find that in almost every case the ranking is preserved across the datasets – and in the case where the rankings switch, the performance differential between the two models is small – suggesting there is little information to be gained by evaluating on an arbitrarily larger dataset. Additionally, we conduct bootstrapping to estimate the variance of model performance on the curated dataset. Concretely, for each task, we sample with replacement 500 times from the relevant question set and evaluate the performance of Gemini-Pro Vision (middle-performing model capable of both tasks) on the sample. Repeating this process 100k times yields a mean accuracy and variance of () and () for the two tasks. This low variance provides further evidence that our curated dataset is sufficiently representative.
We also analyse the degree to which the noisy data can be used to provide few-shot examples. We experimented with both ‘random’ and ‘similar’ (selected based on similarity) examples. The results for Gemini 1.5 Flash are shown in Tab.8. We find that model performance is sensitive to the prompt and with some prompt structures, the presence of examples decreases performance. However, with certain prompts as the results show, incorporating potentially noisy examples can quantitatively improve performance compared to the 0-shot setting. Additionally, a qualitative review of the outputs suggests that the model’s responses more closely follow the instructed format when examples are included, which reduces the need for an LLM to parse the correct answer from the LMM output. The finding that performance can decline with increased shots in certain scenarios is consistent with other works, e.g. [25].
4.5 Error analysis
Fig.8 displays a decomposition of performance by arXiv category for the FigureCaption task for a subset of the evaluated models. We find that the relative rankings of models remain broadly consistent across categories. However, there are clear differences in performance across the categories with most models scoring highly on ‘stat’ questions and much lower on ‘q-bio’. Refusal rates to provide a multiple-choice answers for every question was low among the evaluated LMMs, though models with a higher proclivity for verbose outputs tended to be less decisive. We include an analysis of performance across different properties of the figure-caption pairs making up each question, including the publication year of the source arXiv paper, caption length and figure size, in Fig.8. In general, there is no clear evidence of publication year and figure size having any impact on model performance, though a slight macro-trend of decreasing performance with increasing caption length is observed.
4.6 Caption generation
We extend the breadth of our evaluation by assessing the capabilities of the LMMs to generate suitable captions for scientific figures. Initially, we construct a test set by randomly sampling 100 figure-caption pairs from the FigureCaption task. We then prompt a set of test LMM baselines to generate captions for each figure and select the GPT-4o, Gemini-1.5 Pro, Claude 3.5 Sonnet [73] models to evaluate the generated captions. These specific models were chosen as evaluators because they are ‘stronger’ than the test models on most benchmarks and are from different model families, reducing potential bias. The generated captions were shuffled with the true caption and passed to the evaluator models to rank. We report the rankings across all samples in Tab.3. The results clearly delineate preference among the test models with the closed-source models outperforming the open-source models (as they do on SciFIBench). Captions generated by all closed-source models are preferred over the true caption. Conversely, the true captions are preferred over all the open-source model captions. Strong agreement is shown between the evaluator models. Further details can be found in the Appendix.
| Model | Ranking (µ ± SD) | |||
| Claude 3.5 Sonnet | GPT-4o | Gemini 1.5 Pro | Combined | |
| GPT-4 Turbo | 2.27 ± 1.35 | 2.13 ± 1.20 | 2.07 ± 1.19 | 2.16 ± 1.12 |
| Gemini 1.5 Flash | 3.09 ± 1.64 | 2.92 ± 1.63 | 2.98 ± 1.61 | 3.00 ± 1.28 |
| Claude 3 Sonnet | 3.77 ± 2.61 | 4.14 ± 2.53 | 3.54 ± 2.31 | 3.82 ± 1.59 |
| Claude 3 Haiku | 4.00 ± 2.11 | 4.37 ± 2.25 | 4.56 ± 1.98 | 4.31 ± 1.47 |
| Gemini 1.0 Pro Vision | 4.72 ± 2.16 | 4.22 ± 1.92 | 4.33 ± 1.87 | 4.42 ± 1.42 |
| True Caption | 6.22 ± 2.03 | 5.72 ± 1.93 | 5.67 ± 2.17 | 5.87 ± 1.45 |
| Qwen VL Chat | 5.86 ± 2.03 | 5.77 ± 1.67 | 6.06 ± 1.73 | 5.90 ± 1.35 |
| TransCore-M | 7.85 ± 1.51 | 8.05 ± 1.37 | 8.09 ± 1.27 | 8.00 ± 1.18 |
| OmniLMM 3b | 8.57 ± 1.27 | 8.75 ± 1.32 | 8.83 ± 1.13 | 8.72 ± 1.12 |
| Yi VL 6b | 8.65 ± 1.53 | 8.93 ± 1.20 | 8.87 ± 1.22 | 8.82 ± 1.15 |